# Hulk: A Universal Knowledge Translator for Human-Centric Tasks

### v4

@@ Abstract
Human-centric perception tasks, e.g., pedestrian detection, skeleton-based action recognition, and pose estimation, have wide industrial applications, such as metaverse and sports analysis. There is a recent surge to develop human-centric foundation models that can benefit a broad range of human-centric perception tasks. While many human-centric foundation models have achieved success, they did not explore 3D and vision-language tasks for human-centric and required task-specific finetuning. These limitations restrict their application to more downstream tasks and situations. To tackle these problems, we present Hulk, the first multimodal human-centric generalist model, capable of addressing 2D vision, 3D vision, skeleton-based, and vision-language tasks without task-specific finetuning. The key to achieving this is condensing various task-specific heads into two general heads, one for discrete representations, e.g., languages, and the other for continuous representations, e.g., location coordinates. The outputs of two heads can be further stacked into four distinct input and output modalities. This uniform representation enables Hulk to treat diverse human-centric tasks as modality translation, integrating knowledge across a wide range of tasks. Comprehensive evaluations of Hulk on 12 benchmarks covering 8 human-centric tasks demonstrate the superiority of our proposed method, achieving state-of-the-art performance in 11 benchmarks. The code is available on https://github.com/OpenGVLab/Hulk.
@@
Figure 1: Left: Our proposed Hulk is a generalist human-centric perceiver that can tackle major 2D vision, 3D vision, skeleton-based, and vision language tasks using a unified framework. Right: Hulk pushes the limits on various human-centric tasks when compared with specialist models, pretraining models, e.g., MotionBERT [1], and existing generalist models, e.g., UniHCP [2].
@@ I Introduction
RECENT years have witnessed increasing research attention to human-centric tasks, e.g., person re-identification [3, 4, 5, 6], pose estimation [7, 8, 9, 10], and human mesh recovery [1, 11], due to their widespread applications, including sports analysis [12, 13, 14], surveillance [15, 16, 17, 18], augmented reality [19, 20, 21], and video production [22, 23, 24]. $*{}^{*}$izhou Wang and Yixuan Wu are contributed equally. 🖂Shixiang Tang is the corresponding author. This work was done when Yixuan Wu, Xun Guo were interns at Shanghai AI Laboratory and Shixiang Tang was an intern at Sensetime Research. Yixuan Wu is with the School of Medicine, Zhejiang University, Hangzhou, 310027, China. Weizhen He is with the College of Electrical Engineering, Zhejiang University, Hangzhou, 310027, China. Xun Guo is with the Department of Automation, University of Science and Technology of China, Hefei, 230052, China. Shixiang Tang is with the School of Electrical and Information Engineering, University of Sydney, NSW, Australia. Feng Zhu and Rui Zhao are with SenseTime Group Limited, China. Jian Wu is with School of Public Health, Zhejiang University, Hangzhou, 310058, China. Yizhou Wang, Lei Bai, Tong He, and Wanli Ouyang are with Shanghai AI Laboratory, Shanghai, 200232, China. While most of the previous methods typically emphasized unique designs tailored for individual tasks, involving heavy work for task-specific design, fine-tuning, and deployment, an emerging trend is to explore learning general representative encoders, also known as human-centric foundation models, that can be fine-tuned or prompt-tuned for different human-centric visual tasks. These models leverage large-scale pretraining on diverse datasets, and the high-related human semantics in different human-centric tasks, achieving remarkable performance across a wide range of tasks, particularly in 2D vision applications such as person re-identification [3, 4, 5, 6], pose estimation [7, 8, 9], and human parsing [25, 26, 27]. Some notable examples of such models include SOLIDER [28], HCMoCo [29], MotionBERT [1], PATH [30], HAP [31], and SMPLer-X [32].
@@
To address the challenges, we present Hulk, a Human-centric universal knowledge translator that can, for the first time, handle the major large-scale 2D vision, 3D vision, skeleton-based and vision-language tasks listed in Table I without task-specific finetuning. Unlike previous methods, Hulk unifies data formats and task modeling by condensing diverse task inputs and outputs into four modalities and treating heterogeneous human-centric tasks as modality translation tasks. First, Hulk condenses diverse task inputs and outputs into only two basic formats, e.g., continuous digits and discrete words, which can be further stacked into four modalities, i.e., Text - discrete natural language tokens; Image - normalized RGB images with continuous values from 0 to 1; Sparse Label - 2D/3D continuous location coordinates plus discrete words as its semantic meanings; Dense Label - per-pixel discrete natural language tokens such as segmentation maps. Compared with existing human-centric models, we avoid designing heterogeneous task heads but explore using four modality tokenizers/de-tokenizers to cover diverse human-centric tasks. These more concise formats facilitate flexibility and scalability for various tasks and datasets. Second, Hulk reformulates diverse human-centric multimodal tasks into the modality translation task and accordingly unifies different task-specific model designs into the encoder-decoder architecture which is similar to models in the machine translation task. The intuition behind this is that if the generalist human-centric perceiver clearly perceives 2D and 3D information of human beings in an image, skeletons, or pedestrian appearance descriptions, we only need to teach it to translate between any input and output modalities. Following recent advances in machine translation [41, 42], Hulk unifies different task-specific designs into five simple and effective components: a modality-specific tokenizer, a modality-shared encoder, a modality-specific indicator to tell the decoder which modality the inputs should be translated to, a modality-shared decoder, and a modality-specific de-tokenizer after the decoder. The tokenizers embed multimodal data into token sequences in a common manifold space. The modality-shared encoder then extracts general human-centric representations from the token sequence. The modality-specific indicator guides the modality-shared decoder to translate the human-centric representations into the tokens of the output modality. Finally, the modality-specific de-tokenizer reconstructs the tokens into the desired output modalities to complete any human-centric task.
@@
  • To the best of our knowledge, Hulk is the first generalist human-centric perceiver achieving competitive results on major large-scale 2D vision, 3D vision, skeleton-based and vison-language multimodal human-centric tasks [^1] Tasks such as person re-identification and person search are not explored due to human ethical considerations., including pedestrian attribute recognition, pedestrian detection, 2D pose estimation, human parsing, pedestrian image caption, skeleton-based action recognition, 3D pose estimation and 3D human mesh recovery.
@@
[…]
@@ II Related Work
II-A Multimodal Foundation Models
@@
Alignment with paired multimodal data, as the common way to develop multimodal foundation models, can be generally divided into three different categories: (1) Contrastive methods, which utilize contrastive loss to pull features with the same content close and push features with different contents apart. A famous example is CLIP [43], which aligns image-text pairs on websites, achieving remarkable zero-shot classification performance. VideoCLIP [44] and AudioCLIP [45] further extend this idea to alignments text with video and audio, respectively. ImageBind [46] binds five different modalities with images, achieving strong zero-/few-shot performance on cross-modality recognition tasks. (2) Multimodel reconstruction methods, which learn modality-agnostic representation through multi-modal masked autoencoding [47, 48]. For instance, MultiMAE [49] utilizes a modality-share encoder with modality-specific decoders to reconstruct masked RGB, depth and semantic patches, achieving desirable performance on downstream tasks in these modalities. MaskedVLM [50] and SMAUG [51] adopt masked image/text modeling to learn general representations for vision-language tasks. (3) Multimodal large language models (MLLM) [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62] proposes a new approach to handling diverse perception tasks: aligning all modalities into language and formatting all perception tasks into Visual Question Answering (VQA) tasks. Leveraging effective adaptation methods, e.g., LoRA [63] and LLAMA-Adapter [64], MLLM generates language answers using the given samples in different modalities, including image [54, 55, 56], video [57, 58], point cloud [59, 58, 65, 66, 60] and audio [58, 61]. Despite their versatility, MLLMs have not attained state-of-the-art performance in some perception tasks [52, 53], such as detection and mesh recovery.
@@
II-B Human-Centric Foundation Models
@@ III Human-centric 2D Vision, 3D Vision, Skeleton-based and Vision-Language Tasks
The inputs and outputs of human-centric tasks can be categorized into 4 different modalities: Images - RGB images, which are represented as matrices with elements normalized to range between 0 and 1; Texts - discrete words; Sparse Labels - the coordinates with its semantic meaning and locations, e.g., “nose, [0.45, 0.26]”; Dense labels - per-pixel semantics represented by discrete language tokens. In the following, we show that eight major human-centric tasks can be viewed as the translations among the four modalities. Examples are illustrated in Table II.
@@ IV Hulk: A Generalist Human-centric
We now introduce Hulk, a general human-centric knowledge translator, which can tackle human-centric 2D vision, 3D vision, skeleton-based, and vision-language tasks in one model without any task-specific adaption. The key challenge is to unify the heterogeneous representations of the inputs and outputs. Our Hulk has two innovations to address the problem. First, our Hulk proposes that heterogeneous inputs and outputs can be unified into four different modalities, i.e., texts, images, sparse labels (locations or location sequences) and dense labels (pixel-wise semantic words), which can be further simplified into two basic input and output formats of discrete semantic words and continuous values such as location coordinates. For instance, the pedestrian caption produces words that describe the appearance, age, dress-ups, etc.; the human parsing task predicts pixel-wise semantic words of the given image; the 3D pose estimation task predicts both the location of joints and their corresponding semantic words such as the nose and the right winkle. Second, our Hulk proposes a modality translation task where the major 2D vision, 3D vision, skeleton-based, and vision-language human-centric tasks can be viewed as its special case. The intuition of the modality translation task is that if a model well understands all human-centric knowledge, it only needs to translate from the input modality to the output modality to handle human-centric tasks, e.g., human mesh recovery task translates the image into sparse labels that are 3D location of mesh vertices, and skeleton-based action recognition task translates the human keypoint temporal sequences into texts that describe the action semantic names.
@@ IV-B Tokenizer and De-Tokenizer
One of the most critical challenges to developing the generalist human-centric perceiver is to tackle the heterogeneity of the inputs and outputs. Unlike previous methods that utilized modality-specific projectors and task-specific heads for generating diverse inputs and outputs, our approach, Hulk, simplifies this process. Hulk employs a streamlined interface involving the tokenization of task inputs and the de-tokenization of representations into task outputs. This is achieved using just two types of blocks: a tokenizer and a de-tokenizer, as depicted in Fig. 3. Specifically, one block is designed to classify semantic texts, and the other to regress locations. This design is based on the fundamental perception task objectives: to recognize semantic meanings and find their locations. We further showcase these two types of blocks can be stacked to construct the tokenizers and de-tokenizers of four modalities in major human-centric tasks.
@@ IV-B2 Stacking Blocks to modality-specific Tokenizers and De-tokenizers
$\hat{\mathbf{y}}_{T}$ $=\mathcal{Q}_{T}(\mathbf{q})=[\mathcal{H}^{s}(\mathbf{q}_{1},\mathbf{v}),\mathcal{H}^{s}(\mathbf{q}_{2},\mathbf{v}),...,\mathcal{H}^{s}(\mathbf{q}_{N},\mathbf{v})],$ (9)
@@
=$=\{\hat{\mathbf{y}}_{j}=\text{argmax}_{k\in[1,C]}\mathbf{v}^{\top}\text{Upsample}(\hat{\mathbf{f_{j}}})\},j\in[1,N^{\prime}],$
@@ V-A1 Training Datasets
Image caption
CUHK-PEDES [90] 12,206,283
@@
3D pose estimation
Human3.6M [71] 1,895,710
@@
Total
42 30,187,836 $\approx$30M

### v5

@@ Abstract
Human-centric perception tasks, e.g., pedestrian detection, skeleton-based action recognition, and pose estimation, have wide industrial applications, such as metaverse and sports analysis. There is a recent surge to develop human-centric foundation models that can benefit a broad range of human-centric perception tasks. While many human-centric foundation models have achieved success, they did not explore 3D and vision-language tasks for human-centric and required task-specific finetuning. These limitations restrict their application to more downstream tasks and situations. To tackle these problems, we present Hulk, the first multimodal human-centric generalist model, capable of addressing 2D vision, 3D vision, skeleton-based, and vision-language tasks without task-specific finetuning. The key to achieving this is condensing various task-specific heads into two general heads, one for discrete representations, e.g., languages, and the other for continuous representations, e.g., location coordinates. The outputs of two heads can be further stacked into four distinct input and output modalities. This uniform representation enables Hulk to treat diverse human-centric tasks as modality translation, integrating knowledge across a wide range of tasks. Comprehensive evaluations of Hulk on 12 benchmarks covering 8 human-centric tasks demonstrate the superiority of our proposed method, achieving state-of-the-art performance in 11 benchmarks. The code will be available on https://github.com/OpenGVLab/Hulk.
@@
Figure 1: Left: Our proposed Hulk is a generalist human-centric perceiver that can tackle several common 2D vision, 3D vision, skeleton-based, and vision language tasks using a unified framework. Right: Hulk pushes the limits on various human-centric tasks when compared with specialist models, pretraining models, e.g., MotionBERT [1], and existing generalist models, e.g., UniHCP [2].
@@ I Introduction
RECENT years have witnessed increasing research attention to human-centric tasks, e.g., person re-identification [3, 4, 5, 6], pose estimation [7, 8, 9, 10], and human mesh recovery [1, 11], due to their widespread applications, including sports analysis [12, 13, 14], surveillance [15, 16, 17, 18], augmented reality [19, 20, 21], and video production [22, 23, 24]. Yizhou Wang and Yixuan Wu are contributed equally. 🖂Shixiang Tang is the corresponding author. This work was done when Yixuan Wu, Xun Guo were interns, Yizhou Wang was a research assistant at Shanghai Artificial Intelligence Laboratory. Yizhou Wang and Wanli Ouyang are with Department of Information Engineering, The Chinese University of Hong Kong, Hong Kong, China. Yixuan Wu is with the School of Medicine, Zhejiang University, Hangzhou, 310027, China. Weizhen He is with the College of Electrical Engineering, Zhejiang University, Hangzhou, 310027, China. Xun Guo is with the Department of Automation, University of Science and Technology of China, Hefei, 230052, China. Shixiang Tang is with CUHK Interdisciplinary Artificial Intelligence Research Institute, Hong Kong, China. Feng Zhu and Rui Zhao are with SenseTime Group Limited, China. Jian Wu is with School of Public Health, Zhejiang University, Hangzhou, 310058, China. Lei Bai, Tong He, and Wanli Ouyang are with Shanghai Artificial Intelligence Laboratory, Shanghai, 200232, China. While most of the previous methods typically emphasized unique designs tailored for individual tasks, involving heavy work for task-specific design, fine-tuning, and deployment, an emerging trend is to explore learning general representative encoders, also known as human-centric foundation models, that can be fine-tuned or prompt-tuned for different human-centric visual tasks. These models leverage large-scale pretraining on diverse datasets, and the high-related human semantics in different human-centric tasks, achieving remarkable performance across a wide range of tasks, particularly in 2D vision applications such as person re-identification [3, 4, 5, 6], pose estimation [7, 8, 9], and human parsing [25, 26, 27]. Some notable examples of such models include SOLIDER [28], HCMoCo [29], MotionBERT [1], PATH [30], HAP [31], and SMPLer-X [32].
@@
To address the challenges, we present Hulk, a Human-centric universal knowledge translator that can, for the first time, handle several common large-scale 2D vision, 3D vision, skeleton-based and vision-language tasks listed in Table I without task-specific finetuning. Unlike previous methods, Hulk unifies data formats and task modeling by condensing diverse task inputs and outputs into four modalities and treating heterogeneous human-centric tasks as modality translation tasks. First, Hulk condenses diverse task inputs and outputs into only two basic formats, e.g., continuous digits and discrete words, which can be further stacked into four modalities, i.e., Text - discrete natural language tokens; Image - normalized RGB images with continuous values from 0 to 1; Sparse Label - 2D/3D continuous location coordinates plus discrete words as its semantic meanings; Dense Label - per-pixel discrete natural language tokens such as segmentation maps. Compared with existing human-centric models, we avoid designing heterogeneous task heads but explore using four modality tokenizers/de-tokenizers to cover diverse human-centric tasks. These more concise formats facilitate flexibility and scalability for various tasks and datasets. Second, Hulk reformulates diverse human-centric multimodal tasks into the modality translation task and accordingly unifies different task-specific model designs into the encoder-decoder architecture which is similar to models in the machine translation task. The intuition behind this is that if the generalist human-centric perceiver clearly perceives 2D and 3D information of human beings in an image, skeletons, or pedestrian appearance descriptions, we only need to teach it to translate between any input and output modalities. Following recent advances in machine translation [41, 42], Hulk unifies different task-specific designs into five simple and effective components: a modality-specific tokenizer, a modality-shared encoder, a modality-specific indicator to tell the decoder which modality the inputs should be translated to, a modality-shared decoder, and a modality-specific de-tokenizer after the decoder. The tokenizers embed multimodal data into token sequences in a common manifold space. The modality-shared encoder then extracts general human-centric representations from the token sequence. The modality-specific indicator guides the modality-shared decoder to translate the human-centric representations into the tokens of the output modality. Finally, the modality-specific de-tokenizer reconstructs the tokens into the desired output modalities to complete any human-centric task.
@@
  • To the best of our knowledge, Hulk is the first generalist human-centric perceiver achieving competitive results on several common large-scale 2D vision, 3D vision, skeleton-based and vison-language multimodal human-centric tasks [^1] Tasks such as person re-identification and person search are not explored due to human ethical considerations., including pedestrian attribute recognition, pedestrian detection, 2D pose estimation, human parsing, pedestrian image caption, skeleton-based action recognition, 3D pose estimation and 3D human mesh recovery.
@@
  • Trained on a massive collection of about 30 million labeled human-centric datasets of eight tasks, Hulk can directly handle a broad range of tasks without further task-specific adaptation, even pushing the performance limits of the existing state-of-the-art methods. Specifically, as shown in Fig. 1, Hulk outperforms current leading human-centric specialist models and pretraining models by +1.5% mIoU on CIHP for human paring, +3.1% mA on RAPv2 for pedestrian attribute recognition, +2.0% AP on AIC for 2D pose estimation, -0.7% MR-2 (Missing rate) on CrowdHuman for pedestrian detection, -1.4 MPVPE (Mean-Per-Vertex-Position-Error) and -2.0 MPJPE (Mean-Per-Joint-Position-Error) on 3DPW for mesh recovery and 3D pose estimation, respectively.
@@ II Related Work
II-A Specialist Models on Human-centric Perception Tasks
Researchers have been developing specialist models tailored for individual tasks. To address challenges in human-centric perception tasks, four key strategies have been employed to design these specialist models: (1) Incorporating multi-granularity feature extraction. Tailored backbones, e.g., HRNet [43] and Swin-Transformer [44] are designed to extract representative human-centric features in 2D pose estimation [45, 43, 46] and pedestrian detection tasks [47], pushing the limits on human-centric benchmarks. (2) Leveraging the inherent relationships among body parts and attributes. For example, GCN-based methods [48, 49, 50, 51, 52, 53, 54] and transformer-based methods [55, 56, 57] have achieved satisfactory results in skeleton-based action recognition tasks, as they can effectively model the positional relationships among human joints. In addition, modeling the co-occurrence of human attributes [58, 59] has also facilitated attribute recognition and human parsing. (3) Designing for complex scenarios. As the information of a pedestrian in the crowd might be hard for models to learn directly, progressive refinement [47] and model distillation [45] are proposed for better performance. (4) Utilizing knowledge from related tasks. For example, ED-Pose [60] and BUCTD [61] utilize detection information to aid 2D pose estimation, while [62] and [63] use pose supervision to enhance human part segmentation. Despite the achievements, specialist models involve different task-specific designs, leading to multiple times deployments for multiple tasks. In this paper, we propose to unify diverse human-centric tasks into modality transition tasks to learn them all at once.
II-B Multimodal Foundation Models
@@
Alignment with paired multimodal data, as the common way to develop multimodal foundation models, can be generally divided into four different categories: (1) Contrastive methods, which utilize contrastive loss to pull features with the same content close and push features with different contents apart. A famous example is CLIP [64], which aligns image-text pairs on websites, achieving remarkable zero-shot classification performance. VideoCLIP [65] and AudioCLIP [66] further extend this idea to alignments text with video and audio, respectively. ImageBind [67] binds five different modalities with images, achieving strong zero-/few-shot performance on cross-modality recognition tasks. (2) Multimodel reconstruction methods, which learn modality-agnostic representation through multi-modal masked autoencoding [68, 69]. For instance, MultiMAE [70] utilizes a modality-share encoder with modality-specific decoders to reconstruct masked RGB, depth and semantic patches, achieving desirable performance on downstream tasks in these modalities. MaskedVLM [71] and SMAUG [72] adopt masked image/text modeling to learn general representations for vision-language tasks. (3) Multimodel image generation methods proposes to deal with diverse perception tasks as generating the target images with prediction results, e.g., red bounding boxes for object detection [73], colorful region masks for instance segmentation [74, 75, 76]. Although these methods unify 2D vision tasks, it is hard to extend them in the field of 3D vision. (4) Multimodal large language models (MLLM) [77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87] proposes a new approach to handling diverse perception tasks: aligning all modalities into language and formatting all perception tasks into Visual Question Answering (VQA) tasks. Leveraging effective adaptation methods, e.g., LoRA [88] and LLAMA-Adapter [89], MLLM generates language answers using the given samples in different modalities, including image [79, 80, 81], video [82, 83], point cloud [84, 83, 90, 91, 85] and audio [83, 86]. Despite their versatility, MLLMs have not attained state-of-the-art performance in some perception tasks [77, 78], such as detection and mesh recovery.
@@
II-C Human-Centric Foundation Models
@@ III Human-centric 2D Vision, 3D Vision, Skeleton-based and Vision-Language Tasks
The inputs and outputs of human-centric tasks can be categorized into 4 different modalities: Images - RGB images, which are represented as matrices with elements normalized to range between 0 and 1; Texts - discrete words; Sparse Labels - the coordinates with its semantic meaning and locations, e.g., “nose, [0.45, 0.26]”; Dense labels - per-pixel semantics represented by discrete language tokens. In the following, we show that eight common human-centric tasks can be viewed as the translations among the four modalities. Examples are illustrated in Table II.
@@ IV Hulk: A Generalist Human-centric
We now introduce Hulk, a general human-centric knowledge translator, which can tackle human-centric 2D vision, 3D vision, skeleton-based, and vision-language tasks in one model without any task-specific adaption. The key challenge is to unify the heterogeneous representations of the inputs and outputs. Our Hulk has two innovations to address the problem. First, our Hulk proposes that heterogeneous inputs and outputs can be unified into four different modalities, i.e., texts, images, sparse labels (locations or location sequences) and dense labels (pixel-wise semantic words), which can be further simplified into two basic input and output formats of discrete semantic words and continuous values such as location coordinates. For instance, the pedestrian caption produces words that describe the appearance, age, dress-ups, etc.; the human parsing task predicts pixel-wise semantic words of the given image; the 3D pose estimation task predicts both the location of joints and their corresponding semantic words such as the nose and the right winkle. Second, our Hulk proposes a modality translation task where several common 2D vision, 3D vision, skeleton-based, and vision-language human-centric tasks can be viewed as its special case. The intuition of the modality translation task is that if a model well understands all human-centric knowledge, it only needs to translate from the input modality to the output modality to handle human-centric tasks, e.g., human mesh recovery task translates the image into sparse labels that are 3D location of mesh vertices, and skeleton-based action recognition task translates the human keypoint temporal sequences into texts that describe the action semantic names.
@@ IV-B Tokenizer and De-Tokenizer
One of the most critical challenges to developing the generalist human-centric perceiver is to tackle the heterogeneity of the inputs and outputs. Unlike previous methods that utilized modality-specific projectors and task-specific heads for generating diverse inputs and outputs, our approach, Hulk, simplifies this process. Hulk employs a streamlined interface involving the tokenization of task inputs and the de-tokenization of representations into task outputs. This is achieved using just two types of blocks: a tokenizer and a de-tokenizer, as depicted in Fig. 3. Specifically, one block is designed to classify semantic texts, and the other to regress locations. This design is based on the fundamental perception task objectives: to recognize semantic meanings and find their locations. We further showcase these two types of blocks can be stacked to construct the tokenizers and de-tokenizers of four modalities in common human-centric tasks.
@@ IV-B2 Stacking Blocks to modality-specific Tokenizers and De-tokenizers
$\hat{\mathbf{y}}_{T}$ $=\mathcal{Q}_{T}(\mathbf{q})=[\mathcal{H}^{s}(\mathbf{q}_{1},\mathbf{v}),\mathcal{H}^{s}(\mathbf{q}_{2},\mathbf{v}),.,\mathcal{H}^{s}(\mathbf{q}_{N},\mathbf{v})],$ (9)
@@
$=\{\hat{\mathbf{y}}_{j}=\text{argmax}_{k\in[1,C]}\mathbf{v}^{\top}\text{Upsample}(\hat{\mathbf{f_{j}}})\},j\in[1,N^{\prime}],$
@@ V-A1 Training Datasets
Image caption CUHK-PEDES [114] 12,206,283
@@
3D pose estimation Human3.6M [96] 1,895,710
@@
Total 42 30,187,836 $\approx$30M
V-A2 Evaluation Datasets
V-B Implementation Details